OpenStudy (anonymous):

tan(arcsin(x/2) please just give me the answer

OpenStudy (anonymous):

opposite=x, hypotenuse=2 2^2=x^2 +(b^2) b=sqrt(4-x^2) tan=opp/adj=x/sqrt(4-x^2)

OpenStudy (anonymous):

would that be the exact answer?